FREE INDUCTION DECAY FROM CARBONYL SULFIDE IN THE THZ SPECTRAL RANGE.

\underline{DAMIEN BIGOURD}, ARNAUD CUISSET, SOPHIE MATTON, FRANCIS HINDLE, ERIC FER-TEIN, ROBIN BOCQUET,WEIDONG CHEN, GAEL MOURET, Laboratoire de Physico-Chimie de l’Atmosphère, CNRS UMR-8101, Université du Littoral Côte d’Opale, 189A Ave. Maurice Schumann, 59140 Dunkerque, France.

A typical pump-probe pulsed THz set-up have been used for Time Domain Spectroscopy experiment in order to investigate free induction decay in carbonyle sulfide (COS).^{a} In the detected time domain signal, a transmitted pulse followed by commensurate transients pulses is observed. The molecules of COS respond to the terahertz excitation by reradiating a Free Induced Decay (FID) pulses : the molecules absorb the incident pulse and then a periodic rephasing and dephasing of the entire 60 excited equally spaced transitions occurs during the FID.^{b} Therefore the sample emits a series of terahertz pulse at the rate equal to the frequency separation between adjacent lines of 2 B, where B is the rotational constant,. The coherent transient pulse are so observe mainly in the case of linear or symmetric molecules. A model^{b} based on Maxwell-Bloch equation has been employed to estimate parameters such as the transition linewidth, the rotational and centrifugal distortion constant with the temporal form.

aY. Cai, I. Brener, J. Federici Appl. Phys. Lett., \underline{77, (2076),1997}

bH. Harde, S. Keiding, D. Grischkowsky Phys. Rev. Lett., \underline{66, (1834),1991}